In one episode of the 1960s television sitcom, Gilligan’s Island, the famous “professor” constructed voltaic cells to use as substitutes for their radio’s dead batteries. A single D-cell has an emf of 1.5 V. Which scraps of metal from their damaged boat, the Minnow, could best be used to create a 1.5 V voltaic cell? Assume that coconuts make great beakers and that seawater is a terrific electrolyte.

Iron/iron(II) (the anchor) -0.44 V

Lead/lead(II) (fishing weights) -0.126 V

Silver/silver(I) (Mrs. Howell’s brooch) -0.799 V

Aluminum/aluminum(III) (the boat’s flagpole) -1.677 V
Answer #1
Standard Potential Chart
Metal/Ion Potential (V)
Fe/Fe2+ -0.44
Pb/Pb2+

-0.126

Ag/Ag+ -0.799
Al/Al3+ -1.677

Now, 2 (Ag+(aq) + e Ag(s)) ; EAg+/Ag = (+)0.799 *2 = (+)1.598 V

Pb2+ + 2e Pb ; EPb2+/Pb = (+) 0.126 V

Since Ag has more reduction tendency than Pb, Ag electrode would be cathode & Pb electrode would be anode.

Cell reaction: 2 Ag+(aq) + Pb(s) 2Ag(s) + Pb2+(aq)

Ecell = Ecathode - Eanode = EAg+/Ag - EPb2+/Pb = 1.598 - 0.126 V = 1.472 V 1.5 V.

Ans: Lead (Fishing weight) & Silver (Mrs. Howell's Brooch) would be suitable to construct a 1.5 V cell.